OpenStudy (anonymous):

Carol conducted a survey of four school cafeterias to find the number of students who like tacos for lunch. The results of her survey are recorded in the table below: School Cafeteria Survey School Total Number of Students in the Cafeteria Number of Students Who Like Tacos A 45 12 B 40 15 C 27 9 D 32 13 Which school has the greatest percentage of students who like tacos for lunch? School A School B School C School D @StudyGurl14

OpenStudy (studygurl14):

School A: 12/45 = ? School B: 15/40 = ? School C: 9/27 = ? School C: 13/32 = ?

OpenStudy (anonymous):

Is it school d? @StudyGurl14

OpenStudy (studygurl14):

yes. good job